Calistoga Produce estimates bad debt expense at 0.5% of credit sales. The company reported accounts receivable and allowance for uncollectible accounts of $471,000 and $1,650, respectively, at December 31, 2017. During 2018, Calistoga's credit sales and collections were $315,000 and $319,000, respectively, and $1,720 in accounts receivable were written off. Calistoga's ending balance in its allowance for uncollectible accounts at December 31, 2018, is: $1,505. $1,575. $1,720. $1,650.
Answer #1
Bad debt expense = 315000*0.5%= 1575
Allowance for uncollectible accounts at December 31, 2018 = 1650-1720+1575= 1505